OpenStudy (toxicsugar22):

can anyone help me with this one P=A/1+RT solve for t

OpenStudy (anonymous):

\[P=\frac{A}{1+RT}\]?

OpenStudy (amistre64):

as is; not accounting for typos: 1+RT = A/P RT = (A/P)-1 T = (A/P)-1 -------- R

OpenStudy (amistre64):

A-P ---- might be the simplified version PR

OpenStudy (anonymous):

damn i wrote it all out and it disappeared!

OpenStudy (toxicsugar22):

so can it be T=A-P/PR

OpenStudy (toxicsugar22):

is that the answer

OpenStudy (anonymous):

\[\frac{A}{P}=1+RT\] \[\frac{A}{P}-1=RT\] \[\frac{A}{RP}-\frac{1}{R}=T\]

OpenStudy (amistre64):

yes

OpenStudy (toxicsugar22):

so can it be t=a-p/pr

OpenStudy (toxicsugar22):

can that be the answer

OpenStudy (amistre64):

the answer can be: a-p --- pr